Two moles HI were heated in a sealed tube at 440C till the equilibrium was reached.HIwas found to be 22% decomposed.The equilibrium constant for dissociation is?

Grade:11

2 Answers

pratham ashish
14 years ago

Hi,

there r 2of eq. constant

K& K

           rxn       2HI -> H2 +I2

  moles initally   2        0     0

moles @eqlb.   2-2x    x      x   

   

                                Kc   =  x*x/{2-2x}^2 = (0.22)^2/4*(0.78)^2 = 0.19888

 

conc.(C)= P/RT  for gas phase rxn 

                                         RT *(x)= PH2 = PI2   = 0.19888        (same as   Kc bcz  power o RT cancel out  )

                            

                                             

 


devesh
3 years ago

there are 2 equilibrium constant  Kc and Kp
The reaction,                  2HIH2+I2
moles initially                     2            0        0
moles of equilibrium       22x     x        x
Kc=(22x)2x×x
=(22x)2x×x=4×(0.78)2(0.22)2
=Concentration(C) =RTP= for gas phase reaction
RT×X=PH2=PI2=0.19888